An investor wishes to realize a return of 12 per cent on a total of 2 investments. If she has ₱10,000.00 invested at 10 per cent, how much additional money should be invested at 16 percent?

-Hello. Sorry for bothering but I really need your help. I'm working at it for hours but I can't really answer it. Please. THANK YOU and MORE POWER!

Answer by richwmiller(17219)

We know 3 rates and one amount invested.
0.1*10000+0.16*x=0.12*(10000+x)
Multiply out.
1000+0.16*x=1200+0.12*x
Combine like terms.
0.04*x=200
Isolate x.
x=200/0.04
Divide.
x=5000
Check.
0.1*10000+0.16(x)=0.12*(10000+x)
1000+0.16*5000=0.12*(10000+5000)
1000+800=0.12*(15000)
1800=1800
Ok